OpenStudy (anonymous):

evaluate 3^-3*3^4*3*3^-5. A.-27 B.1/27 C.-81 D.1/81

OpenStudy (anonymous):

@AnswerMyQuestions @iGreen @Samus_Aran @samanthagreer

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@Destinymasha @iGreen

OpenStudy (anonymous):

Sorry, dont remember how to do these

OpenStudy (igreen):

\(3^{-3} \times 3^4 \times 3 \times 3^{-5}\) Add all the exponents and keep the base: \(3^{-3 + 4 + 1 + (-5)} \rightarrow~?\)

OpenStudy (anonymous):

umm

OpenStudy (igreen):

Just add -3 + 4 + 1 + -5

OpenStudy (anonymous):

-3

OpenStudy (igreen):

Yes, so we have: \(3^{-3}\) Can you solve that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-27

OpenStudy (igreen):

Oh, well we can change this into a fraction using the rule: \(x^{-1} \rightarrow \dfrac{1}{x^{1}}\) So we have: \(\dfrac{1}{3^3}\) What's 3 * 3 * 3?

OpenStudy (anonymous):

27

OpenStudy (anonymous):

so 1/27

OpenStudy (igreen):

Yep, you got it.

OpenStudy (anonymous):

yes thx! i will tagyou in more